Complete { (2, 0,3) } to form an orthogonal basis of R^3.



1
Expert's answer
2021-07-07T15:14:50-0400

Let v = <2,0,3>

We need to find three vectors in R3 such that they form an orthogonal basis.


Let the vectors be u1, u, u.


Let v = "\dfrac{1}{3}" <"\dfrac{2}{3}, 0, \dfrac{3}{3}" >


v "\dfrac{1}{3}" u1



Hence, we compute that u1 = <"\dfrac{2}{3}, 0, \dfrac{3}{3}" >


Our first goal is to find the vectors u2 and u3  such that {u,u2 , u3} is an orthogonal basis for R3 .


Let there be a vector x = <x, y, z>

such that x . u1 = 0.


Hence form the above equation we have


"\dfrac{2}{3} x + z = 0"


Now let vector u2 = <-3, 0, 2> satisfies the above relation.


So, we conclude that u2 = <-3, 0, 2>



We can find u3 = u1 X u2

u3 = <"\dfrac{2}{3}, 0, \dfrac{3}{3}" > X < - 3, 0, 2>


u"\begin{vmatrix}
 i & j &k\\
 2/3 & 0 & 1\\
-3 & 0 &2
\end{vmatrix}"


u3 = <0, "-\dfrac{13}{3}" , 0>



Thus we have found the vectors u1, u2 and u.

To normalize them we divide the vectors by their length. Let v1 , vand v3 be the corresponding unit vectors.



v"\dfrac{3}{\sqrt{13}} <\dfrac{2}{3}, 0, 1>"


v2 = "\dfrac{1}{{\sqrt{13}}} <-3, 0, 2>"


v"\dfrac{3}{13} <0, -\dfrac{13}{3}, 0>"


{v1, v2, v3} = {"\dfrac{3}{\sqrt{13}} <\dfrac{2}{3}, 0, 1>" , "\dfrac{1}{{\sqrt{13}}} <-3, 0, 2>" , "\dfrac{3}{13} <0, -\dfrac{13}{3}, 0>" } form an orthonormal


basis on Rcontaining the vector <2, 0, 3> .